Prove that if f is continuous, then:


∫ (x>0) f(u)(x-u).du= ∫ (x>0) ( ∫ (u>0) f(t) dt) du.





1
Expert's answer
2021-11-04T09:20:38-0400

let


"g(x)=\displaystyle{\int ^x_0}f(u)(x-u)du-\displaystyle{\int ^x_0}(\displaystyle{\int ^u_0}f(t)dt)du"


By the Fundamental Theorem of Calculus:


"\frac{d}{dx}(\displaystyle{\int ^x_0}uf(u)du)=xf(x)"


"\frac{d}{dx}(\displaystyle{\int ^x_0}\displaystyle{\int ^u_0}f(t)dtdu)=\displaystyle{\int ^x_0}f(t)dt"


Notice here we needed both that f and the indefinite integral of f are continuous functions. Now,


"g'(x)=\displaystyle{\int ^x_0}f(t)dt+xf(x)-xf(x)-\displaystyle{\int ^x_0}f(u)du=0"


So by the Mean Value Theorem, g is a constant. Further, observe that 


"g(0)=\displaystyle{\int ^0_0}f(u)(x-u)du-\displaystyle{\int ^0_0}(\displaystyle{\int ^u_0}f(t)dt)du=0-0=0"


Thus, g(x) = 0 for all x. It follows that


"\displaystyle{\int ^x_0}f(u)(x-u)du=\displaystyle{\int ^x_0}(\displaystyle{\int ^u_0}f(t)dt)du"